ECH 4604: Final Exam Fall 2005

Develop a CHEMCAD module in Version 5.5 for the unit operation assigned to you in the table below. The format of the module should be along the lines of the modules given in the class. You can use the existing modules as a template for your module.

Name  Unit Operation
Bailey, Natalee  How to install CHEMCAD
Beard, Michael  Mixer
Blue, Kacie  CSTR using reaction kinetics
Boyd, Rogelio  Heat Exchanger
Brewer, Jared  Splitter
Cardoso, Carlos  PFR using reaction kinetics
Castro, Adirene  Tray Absorber
Charlton, Jacqueline Distillation cost estimation
Chou, Hsiang-Yu  Flash
Cooksey, Shanann  Distillation: Short-cut calculations
Dasilva, Marilia  CSTR using Gibbs reactor
Dekloet, Annette  Packed Absorber
Dotson, Steven  VLE data
Gaffney, Ashley  Batch Reactor
Helkin, Jessica  Distillation: Tray-by-tray calculations
Lawson, Jeffrey  Flash
Lescher, Mark  Tray Absorber
Loshin, Aaron  CSTR using stoichiometry
Mitchell, Jameer  Piping network
Nnaji, Ninika  Tray Absorber
Poplin, Micah  Mixer-Settler
Richardson, Chanthea Packed Absorber
Sellgren, Katelyn  Heat Exchanger
Stjohn, Robert  Flowsheet with recycle
Thompson, Craig  Batch Reactor
Thompson, Kimberly Liquid-Liquid extraction
Togun, Olusegun  Piping network
Valencia, Alejandro  Flowsheet with recycle
Young, Brandon  PFR using reaction kinetics

• This final is due by Thursday, December 12, 2005 at 12:00 noon (this is the scheduled date and time for the final for this course). The module should be written as a Word document and emailed to me at palanki@eng.fsu.edu.

• This exam is worth 2 points. If you do not turn in your module by the specified date and time, you will get a zero for this exam.

• This is not a team assignment; each student is expected to work on this problem individually and the usual honor code rules for exams apply to this final.